OpenStudy (anonymous):

How many different 12-member juries can be made from 150 people?

OpenStudy (anonymous):

150C12 which is also:\[\left(\begin{matrix}150 \\ 12\end{matrix}\right)\]or:\[\frac{ 150! }{ 12! \times 138! }\]

OpenStudy (anonymous):

n! means: n(n - 1)(n - 2)(n - 3) . . . (3)(2)(1)

OpenStudy (anonymous):

172,420,656,389,440,550?

OpenStudy (anonymous):

Here's a shortcut: 150!/[(12!)(138!)] = (150 x 149 x 148 x ... x 141 x 140 x 139 x 138!) / [(12!)(138!)] The "138!" factors will cancel.

OpenStudy (anonymous):

I checked out your answer and it's right.